[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Oct 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Oct 27 20:54:03 UTC 2017


HSU Shift Log for Fri Oct 27 2017 16:53:55 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: The issue with the location of data being plotted for buoy  
42057 was resolved by NDBC.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: New G-16 meso request for 23z-09z. Will need to  
updated late this evening or overnight in order to extend it into Saturday.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171027/3aa938f9/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list